OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x − 3 y = 5x − 1 y = 5x + 3 y = 5x − 19

OpenStudy (karatechopper):

So to do this problem you need to know what Slope-intercept form is. Slope intercept form is: y=mx+b m= Slope b= y-intercept y= y-coordinate x=x coordinate

OpenStudy (anonymous):

d. y=5x-19 y=1 x=4 1= 5(4)-19 1=1

OpenStudy (karatechopper):

Now all you have to do is plug in the slope and the (x,y) coordinate to find the y-intercept ->b

OpenStudy (karatechopper):

er..I guess you got help up there so I am just gonna..let it off now lol.